# Application of AI to formal methods — an analysis of current trends

### v1

@@ Abstract
With artificial intelligence (AI) being well established within the daily lives of research communities, we turn our gaze toward an application area that appears intuitively unsuited for probabilistic decision-making: the area of formal methods (FM). FM aim to provide sound and understandable reasoning about problems in computer science, which seemingly collides with the black-box nature that inhibits many AI approaches. However, many researchers have crossed this gap and applied AI techniques to enhance FM approaches. As this dichotomy of FM and AI sparked our interest, we conducted a systematic mapping study to map the current landscape of research publications. In this study, we investigate the previous five years of applied AI to FM (2019–2023), as these correspond to periods of high activity. This investigation results in 189 entries, which we explore in more detail to find current trends, highlight research gaps, and give suggestions for future research.
@@
journal: Information and Software Technology
[…]
[1]organization=Johannes Kepler University Linz, organization=Institute of Software System Engineering, city=Linz, postcode=4040, country=Austria
[…]
[2]organization=Heinrich Heine University Düsseldorf, organization=Faculty of Mathematics and Natural Sciences, organization=Institute of Computer Science, city=Düsseldorf, postcode=40225, country=Germany
{highlights}
Systematic mapping of AI application within the formal methods domain.
Analyzed tally of 189 primary studies from the last five years (2019–2023).
Data set of all 457 found publications, including the mapped tally of 189 processed studies, made publicly available at https://github.com/hhu-stups/ai4fm-studies.
Formulation of research questions to guide future research.
@@ 1 Introduction
In this study, we ask whether this trend is also observable within the formal methods communities. Formal methods (FM) correspond to a mathematically rigorous approach to software and systems development in which the concern is to provide a formal representation of a hardware, software, or systems engineering problem [7]. This representation then lends itself to mathematical assessments such as correctness, soundness, or well-definedness. Furthermore, these assessments can be easily reproduced and followed by the practitioner. The needed guarantees and the rigorousness of these methods seem to stand in tension with AI applications, which often show nondeterministic behavior and mostly correspond to black-box techniques. However, much like in software development, AI-driven tools might still supplement the formal development process without impeding the correctness of the results. And from personal experience, we noticed an increase in publications applying AI techniques to FM in recent years, suggesting that the FM field is following the overall trend as well. To substantiate this and to obtain an overview of which FM subdomains already utilize AI, we performed this systematic mapping study (SMS) [8]. The primary goal is to observe the development of AI applications quantitatively to the field of FM.
@@
In this work, we report on our journey and the results of our SMS, accessing the quantitative nature of the topic of AI application in the FM domain. This represents a typical goal for this kind of study, as pointed out by Kitchenham et al. [9]. Mainly, we are interested in which FM domains are targeted by AI improvement already, which AI or ML applications have been investigated in the literature so far, where potential research gaps exist, and whether there appears to be consensus within the FM domains of which AI techniques seem most beneficial. A challenge to this assessment is the wide variety of topics on both sides. AI and FM are huge topics, and their potential overlap may be considerable. This study aims to be the first assessment of the general research landscape, i.e., we assume a mainly quantitative lens on the topic. Consequently, we highlight the areas which received more attention in the literature as well as those that received less attention. We refrain, however, from doing a qualitative assessment (e.g., a systematic literature review) our goal is to gain an initial overview of the field’s landscape.
@@
While our search process resulted in a data set of 457 highly relevant publications covering more than 50 years of research applying AI to FM, we investigated only articles from the last five years in more detail. This period is chosen as it corresponds to a peak in the number of publications as indicated by previously mentioned studies and also allows us to focus on the most recent developments. Consequently, we end up with a tally of 191 primary studies, two of which are pure data set presentations. We analyzed the remaining 189 studies to determine which AI techniques were found to be applicable in which FM areas. Further, we derive research suggestions from our insights by pointing out research gaps that might further mature the field. Both the data set of the last five years as well as the complete data set of 457 publications are made publicly available to interested researchers at https://github.com/hhu-stups/ai4fm-studies.
@@ 2.1 Systematic mapping studies
Systematic mapping studies, as proposed by Kitchenham et al. [10], aim to assess the quantitative nature of a research field. Here, the goal is to give an overview of the available research. SMS complement systematic literature reviews by comparing research contributions and their qualitative impact.
SMS are typically performed using a multi-step approach. First, the field of interest is selected. Then, research questions are formulated, and inclusion criteria (IC) and exclusion criteria (EC) are defined. IC and EC will serve as filters to help select relevant results.
After that, the search query is crafted. The goal is to apply the search query to the scientific meta-search engines to retrieve a good corpus of contributions that answer the research questions. With the retrieved corpus, the IC and EC are applied. If indicated, snowballing is done, usually until a stable closure of the corpus is reached. With a corpus constructed, the aim is to answer the research questions.
@@ 2.2 Formal methods
Formal methods (FM) describe a mathematically rigorous approach to design software as well as hardware systems, and are concerned with analysis, validation, and verification at any part of the respective system’s life cycle [11, 12]. Employing FM allows to formulate precise statements of desired functionality or requirements in form of a formal model (or formal specification) while not constraining the possible implementation thereof. The formal model is represented in a mathematically approachable form which lends itself to reasoning and hence allows rigorous analysis of critical properties such as correctness, safety, soundness, or well-definedness.
In the following, we overview FM creation and reasoning methods most relevant to this works.
@@
Model checking [13, 14] explores a program’s or systems’ state space. The state space sets all possible value constellations the program can achieve. In its most basic form, explicit state model checking, model checking aims to explore all reachable states and check if they are faulty, i.e., if they violate any specified properties. If this is the case, a counter-example is found. Otherwise, the system works correctly, as no faulty states are reachable.
There are many flavors besides explicit state model checking. Noteworthy for this study are linear temporal logic (LTL) model checking, symbolic model checking, and statistical model checking. LTL model checking encodes temporal properties of interest into a formula that is checked against execution traces of the state space. Symbolic model checking abstracts the state space and makes a symbolic evaluation of a complex expression, resulting in the ability to handle even infinite state spaces for the price of precision. Statistical model checking investigates state spaces created by assigning probabilities on transitions between states and provide probabilistic guarantees.
@@
Satisfiability modulo theories (SMT) [18] is a more generalized form of the SAT which enriches the problem setting with various theories such as arithmetics, data structures, or set theory.
@@
Under the term of synthesis, we group all attempts automatically or semi-automatically create (parts of) formal models during the formal development process, such as the generation of models from natural language specification or vice versa. We further group the idea of automated model repair under this aspect, where for a model with some violation, a fix is synthesized that rectifies the violation.
@@
In the scope of this mapping study, we may encounter FM approaches and techniques, which do not belong to any previous group. We will group these as other FM techniques. For instance, this includes meta-approaches such as selecting the right verification algorithm, general program analysis, or termination analysis.
@@ 2.3 Artificial intelligence
Artificial Intelligence (AI) is a catch-all phrase for various techniques and approaches that aim to imitate intelligent decision-making. Simmons and Chappell [23] define the term as “[…] behavior of a machine which, if a human behaves in the same way, is considered intelligent”. They further emphasize the problem with the term itself, pointing out the difficulty of precisely defining what it means for a human to be intelligent. Nevertheless, AI is regularly used in daily life and the research community.
An important AI paradigm is machine learning (ML). The goal of ML algorithms is to observe and extract patterns within the data which can then be applied to the automation of a given task  [24]. Instead of manually defining a set of rules for the pattern recognition, machine learning algorithms uncover such rules in an automated fashion, a process referred to as learning. ML algorithms experience the given data in some manner and improve their exhibited performance to solve the task at hand gradually, the more experience they gather.
Below, we list and briefly explain AI and ML algorithms which are most relevant to this mapping study. For the reader familiar with AI, the selection seems mostly ML rather than AI. Indeed, we aimed to find applications of AI to FM, and the search query we will discuss later was prepared accordingly. However, in order to bring a result up front, most contributions use some form of ML.
@@
Neural networks (NNs) [25] build the foundation of what is known today as the area of deep learning [26]. A deep neural network (DNN) consists of a layer of inputs, a layer of outputs, and one or more hidden layers. Inputs are forwarded layer-wise and combined first by a weighted sum and a follow-up non-linear transformation. The clue is the training of a DNN, which starts the network with randomly initialized weights for the summations above but adjusts them gradually in the learning process to converge to the desired function.
@@
Reinforcement learning (RL) [32] is foremost a machine learning paradigm. Given an environment, a set of actions that change the environment, and a reward function that quantifies a given environmental state, the RL agent’s goal is to learn a policy over their available actions which maximizes the cumulative reward over time. Initially, the agent does not know the environment or the task it is supposed to solve but can only learn from feedback through the received rewards [33].
@@
One of the most recent developments is large language models (LLMs) [36] which gained significant popularity and even worldwide attention with the release of ChatGPT [37].
@@
Genetic and evolutionary algorithms (EAs) [38] describe a family of optimization algorithms that are rooted in the phenomenon of evolution. An EA starts with a randomly assembled population of candidates for a given optimization problem and a fitness function. The fitness function measures how well a candidate solves the problem. Parent candidates are chosen via a random selection process that is nonetheless influenced by each individual’s fitness. These are adapted using mutation or recombination to produce a new child population. Repeating this process converges the individuals toward suitable solutions for the problem at hand.
@@
In contrast to the abovementioned techniques, we consider techniques from more classical, statistical ML. These differ from the above methods as they do not depend on deep learning and are typically faster to compute in comparison. This does, however, not imply that they are not competitive. The considered statistical approaches include support vector machines (SVM) [39], logistic and linear regression (LR) [40], k nearest neighbors (KNN) [41], support vector machines (SVM) [42], decision trees (DT) [43] and random forests (RF) [44], or Bayesian inference approaches [45].
@@ 2.4 Contribution classification
We classified the contributions that we selected via the PICO [10] methodology. Thus, we facilitate a clear differentiation between them and allow for a more refined analysis.
The different types of contributions are distinguished as follows:
@@
Contributions that propose an overall style or idea to overcome a problem. An approach is a generalized concept that describes how to react or overcome a problem in depth. The approach remains more theoretical and does not involve tested or empirically proven steps.
@@
Contributions applying an approach, a methodology, or a framework to a real-world problem.
@@
Contributions that very briefly overview how a problem may be solved. Compared to an approach, it does not lay out larger theories but serves as a quick pitch.
@@
An openly accessible data set that claims to be reusable outside of replication of the contribution it was originally used in.
We further divide the publication venues into the following categories:
@@
3 Goal & Research questions
To get an overview of the field, we divide our research questions (RQs) into two main areas. First, RQs 1.1, 1.2 and 1.3 focus on the overall demographics of the research field. With these questions, we want a basic quantitative overview to explore the field’s maturity. Second, RQs 2.1, 2.2, 2.3 and 2.4 focus on the content of the contributions and aim to map out the field, highlighting potential gaps and showcasing quantitative interest in respective sub-topics. Our research questions are defined as follows.
@@ 1. RQ 1:
      Rational: We are interested to know whether the field experiences a growth or decline in interest, or if no trends are detectable at all.
@@
      Rational: We want to explore the field’s maturity, whereby books, book chapters, and journal articles indicate more mature results.
@@
      What are the provided main contributions by the primary studies?
      Rational: We want to learn where the community focuses its efforts, e.g., more on the theoretical groundwork or more on practical application.
@@ 2. RQ 2:
      Rational: We want to know if there is an indication that some AI techniques are more relevant (or at least popular) than others.
@@
      What are the application areas of AI in FM? Are there any commonly found FM techniques or tools? Rational: We are curious about which FM techniques currently get the most attention or if there are dark spots in the literature.
@@
      Rational: Here we want to overlap the results of RQs 2.1 and $2.$
@@
      Are the studies’ employed data sets publicly available? Rational: Especially in machine learning, having the source material available and accessible for experiments benefits the scientific community.
@@ 4 Search strategy
In the following, we explain our approach for aggregating the contributions for this mapping study which we have visualized in Fig. 1. For this, we follow the suggested multi-step approach by Petersen et al. [8], which consists of an initial search in meta-search engines, followed up by a snowballing process. As it will show, our selection of topics provided a challenging environment which has overlapping search terms with related research areas. To avoid a result set of tens of thousands of potential studies, we tweaked our search strategy to deal with these challenges.
A vital role played the IC & EC, which we applied twice. The first application was made after the initial search to remove unrelated contributions that would introduce an overheat into the snowballing. The second time was again after the snowballing to filter wrongly selected contributions.

### v2

@@ Abstract
Context: With artificial intelligence (AI) being well established within the daily lives of research communities, we turn our gaze toward formal methods (FM). FM aim to provide sound and verifiable reasoning about problems in computer science.
Objective: We conduct a systematic mapping study to overview the current landscape of research publications that apply AI to FM. We aim to identify how FM can benefit from AI techniques and highlight areas for further research. Our focus lies on the previous five years (2019–2023) of research.
Method: Following the proposed guidelines for systematic mapping studies, we searched for relevant publications in four major databases, defined inclusion and exclusion criteria, and applied extensive snowballing to uncover potential additional sources.
Results: This investigation results in 189 entries which we explored to find current trends and highlight research gaps. We find a strong focus on AI in the area of theorem proving while other subfields of FM are less represented.
Conclusions: The mapping study provides a quantitative overview of the modern state of AI application in FM. The current trend of the field is yet to mature. Many primary studies focus on practical application, yet we identify a lack of theoretical groundwork, standard benchmarks, or case studies. Further, we identify issues regarding shared training data sets and standard benchmarks.
@@ 1 Introduction
This study asks whether this trend is also observable within the formal methods communities. Formal methods (FM) correspond to a mathematically rigorous approach to software and systems development in which the concern is to provide a formal representation of hardware, software, or systems engineering problem [7]. This representation then lends itself to mathematical assessments such as correctness, soundness, or well-definedness. These assessments can be easily reproduced and followed by the practitioner.
However, the needed guarantees and the rigorousness of these methods seem to conflict with AI applications, which often show nondeterministic behavior and mostly correspond to black-box techniques. The trade-off appears obvious: Injecting AI’s nondeterminism and unpredictability into FM’s rigorousness could lead to faster results and proofs with less human intervention at the price of fundamental guarantees. This point of tension seems interesting and needs further investigation, but it is not the only way how AI can enrich FM. Much like in software development, AI-driven tools might still supplement the formal development process without impeding the correctness of the results. From personal experience, we noticed increased publications applying AI techniques to FM in recent years, suggesting that the FM field also follows the overall trend.
To substantiate this and to obtain an overview of which FM subdomains already utilize AI, we performed this systematic mapping study (SMS) [8]. The primary goal is to observe the development of AI applications quantitatively to the field of FM.
@@
In this work, we report on our journey and the results of our SMS, accessing the quantitative nature of the topic of AI applications in the FM domain. This represents a typical goal for this kind of study, as pointed out by Kitchenham et al. [9]. Mainly, we are interested in which FM domains are targeted by AI improvement already, which AI or ML applications have been investigated in the literature so far, where potential research gaps exist, and whether there appears to be consensus within the FM domains of which AI techniques seem most beneficial. A challenge in this assessment is the wide variety of topics on both sides. AI and FM are huge topics, and their potential overlap may be considerable. This study aims to be the first assessment of the general research landscape, i.e., we assume a mainly quantitative lens on the topic. Consequently, we highlight areas that received more attention in the literature, as well as those that received less attention. However, we refrain from doing a qualitative assessment, e.g., a systematic review of the literature. Our goal is to gain an initial overview of the field’s landscape.
@@
While our search process resulted in a data set of 457 highly relevant publications covering more than 50 years of research applying AI to FM, we investigated only articles from the last five years in more detail. This period is chosen as it corresponds to a peak in the number of publications as indicated by previously mentioned studies and also allows us to focus on the most recent developments. Consequently, we have 191 primary studies, two of which are pure data set presentations. We analyzed the remaining 189 studies to determine which AI techniques were found to be applicable in which FM areas. Further, we derive research suggestions from our insights by pointing out research gaps that might further mature the field. Both the data set of the last five years and the complete data set of 457 publications are made publicly available to interested researchers at https://github.com/hhu-stups/ai4fm-studies.
@@ 2.1 Systematic mapping studies
Systematic mapping studies (SMS), as proposed by Kitchenham et al. [10], aim to assess the quantitative nature of a research field. Here, the goal is to give an overview of the available research. The counterpart to an SMS would be a systematic literature review where the main goal is to assess the quality of contributions.
SMS are typically performed using a multi-step approach. First, the field of interest is selected. Then, research questions are formulated, and inclusion criteria (IC) and exclusion criteria (EC) are defined. IC and EC serve as filters over the collected primary studies to help select relevant results.
After that, the search query is crafted. The goal is to apply the search query to the scientific meta-search engines to retrieve a good corpus of primary studies that answer the research questions. With the retrieved corpus, the IC and EC are applied. If indicated, snowballing is done, usually until a stable closure of the corpus is reached. With a corpus constructed, the aim is to answer the research questions.
@@ 2.2 Formal methods
Formal methods (FM) describe a mathematically rigorous approach to design and assess software as well as hardware systems, and are concerned with analysis, validation, and verification at any part of the respective system’s life cycle [11, 12]. Employing FM allows to formulate precise statements of desired functionality or requirements in form of a formal model (or formal specification) while not constraining the possible implementation thereof. The formal model is represented in a mathematically approachable form which lends itself to reasoning and hence allows rigorous analysis of critical properties such as correctness, safety, soundness, or well-definedness. The following section overviews FM’s creation and this work’s most relevant reasoning methods.
@@
Model checking [13, 14] explores a program’s or system’s state space. The state space sets all possible value constellations that the program can achieve. In its most basic form, model checking aims to explore all reachable states and check if they are faulty, i.e., violate any specified properties. If this is the case, a counterexample is found. Otherwise, the system works correctly, as no faulty states are reachable.
There are different types of model checking, besides classical, explicit state model checking. Noteworthy for this study are linear temporal logic (LTL) model checking, symbolic model checking, and statistical model checking. While explicitly visiting states, LTL model checking focuses on temporal properties encoded in LTL formulas, which are checked against execution traces of the state space. Symbolic model checking abstracts the state space and makes a symbolic evaluation of a complex expression; thus, it shrinks the overall state space for the price of precision. With enough abstraction, it can handle even infinite state spaces for the price of precision. Statistical model checking investigates state spaces created by assigning probabilities on transitions between states and provide probabilistic guarantees.
@@
Satisfiability modulo theories (SMT) [18] is a more generalized form of the SAT, which enriches the problem setting with various theories such as arithmetic, data structures, or set theory.
@@
Under the term of synthesis, we group all attempts automatically or semi-automatically create (parts of) formal models during the formal development process, such as the generation of models from natural language specification or vice versa. We further group the idea of automated model repair under this aspect, where, for a model with some violation, a fix is synthesized that rectifies the violation.
@@
In the scope of this mapping study, we may encounter FM approaches and techniques that do not belong to any previous group. We will group these as other FM techniques. For instance, this includes meta-approaches such as selecting the right verification algorithm, general program analysis, or termination analysis.
@@ 2.3 Artificial intelligence
Artificial Intelligence (AI) is a catch-all phrase for various techniques and approaches that aim to imitate seemingly intelligent behavior. Simmons and Chappell [23] define the term as “[…] behavior of a machine which, if a human behaves in the same way, is considered intelligent”. They further emphasize the problem with the term itself, pointing out the difficulty of precisely defining what it means for a human to be intelligent. Nevertheless, AI is regularly used in daily life and the research community.
An important AI paradigm is machine learning (ML). The goal of ML algorithms is to observe and extract patterns within the data which can then be applied to the automation of a given task  [24]. Instead of manually defining a set of rules for pattern recognition, machine learning algorithms uncover such rules automatically, a process referred to as learning. ML algorithms experience the given data in some manner and improve their exhibited performance to solve the task at hand gradually, the more experience they gather.
Below, we list and briefly explain AI and ML algorithms which are most relevant to this mapping study. For the reader familiar with AI, the selection seems mostly ML rather than AI. Indeed, we aimed to find applications of AI to FM, and the search query we will discuss later was prepared accordingly. However, to bring a result up front, most contributions use some form of ML.
@@
Neural networks (NNs) [25] build the foundation of what is known today as the area of deep learning [26]. A deep neural network (DNN) consists of a layer of inputs, a layer of outputs, and one or more hidden layers. Inputs are forwarded layer-wise and combined first by a weighted sum and a follow-up non-linear transformation. The clue is the training of a DNN, which starts the network with randomly initialized weights for the summations above, but adjusts them gradually in the learning process to converge to the desired function.
@@
Reinforcement learning (RL) [32] is foremost a machine learning paradigm. Given an environment, a set of actions that change the environment, and a reward function that quantifies a given environmental state, the RL agent aims to learn a policy over their available actions that maximizes the cumulative reward over time. Initially, the agent does not know the environment or the task it is supposed to solve, but can only learn from feedback through the received rewards [33].
@@
One of the most recent developments is large language models (LLMs) [36] which gained significant popularity and even worldwide attention with the release of ChatGPT [37]. For the sake of compactness, we count contributions with LLMs towards NLP.
@@
Genetic and evolutionary algorithms (EAs) [38] describe a family of optimization algorithms rooted in evolution. An EA starts with a randomly assembled population of candidates for a given optimization problem and a fitness function. The fitness function measures how well a candidate solves the problem. Parent candidates are chosen via a random selection process that is nonetheless influenced by each individual’s fitness. These are adapted using mutation or recombination to produce a new child population. Repeating this process converges the individuals toward suitable solutions for the problem at hand.
@@
In contrast to the abovementioned techniques, we consider techniques from more classical, statistical ML. These differ from the above methods as they do not depend on deep learning and are typically faster to compute in comparison. This does, however, not imply that they are not competitive. The considered statistical approaches include support vector machines (SVM) [39], logistic and linear regression (LR) [40], k nearest neighbors (KNN) [41], decision trees (DT) [42] and random forests (RF) [43], or Bayesian inference approaches [44].
@@ 2.4 Contribution classification
We classified the contributions into various types. The aim is to show the different approaches to target the same research area. The different types of contributions are distinguished as follows:
@@
Contributions that propose an overall style or idea to overcome a problem. An approach is a generalized concept that details overcoming and resolving a problem. The approach remains more theoretical and does not involve tested or empirically proven steps.
@@
Contributions apply an approach, a methodology, or a framework to a real-world problem.
@@
Contributions that focus on a brief overview or discussion of an idea to solve a problem. Compared to an approach, it does not lay out larger theories but serves as a quick pitch.
@@
An openly accessible data set that claims to be reusable outside of the replication of the contribution it was originally used in.
2.5 Publication venue classification
We further divide the publication venues into the following categories to better assess the respective maturity of the contributions:
@@ 2.5 Publication venue classification
3 Research questions
To get an overview of the field, we divide our research questions (RQs) into two main areas. First, RQs 1.1, 1.2 and 1.3 focus on the overall demographics of the research field. With these questions, we want a basic quantitative overview to explore the field’s maturity. Second, RQs 2.1, 2.2, 2.3 and 2.4 focus on the content of the contributions and aim to map out the field, highlighting potential gaps and showcasing quantitative interest in respective subtopics. Our research questions are defined as follows.
@@ 1. RQ 1:
      Rationale: We are interested in knowing whether the field experiences a growth or decline in interest, or if no trends are detectable at all.
@@
      Rationale: We want to explore the field’s maturity, whereby books, book chapters, and journal articles indicate more mature results.
@@
      What are the main contributions provided by the primary studies?
      Rationale: We want to learn where the community focuses its efforts, e.g., more on the theoretical groundwork or practical application.
@@ 2. RQ 2:
      Rationale: We want to know if there is an indication that some AI techniques are more relevant (or at least popular) than others.
@@
      What are the application areas of AI in FM? Are there any commonly found FM techniques or tools? Rationale: We are curious about which FM techniques currently get the most attention or if there are dark spots in the literature.
@@
      Rationale: Here we want to overlap the results of RQs 2.1 and $2.$
@@
      Are the studies’ employed data sets publicly available? Rationale: Especially in machine learning, having the source material available and accessible for experiments benefits the scientific community.
@@ 4 Search strategy
In the following, we explain our approach for aggregating the contributions for this mapping study which we have visualized in Fig. 1. For this, we follow the suggested multi-step approach by Petersen et al. [8], which consists of an initial search in meta-search engines, followed by a snowballing process. As our selection of topics will show, it provided a challenging environment with overlapping search terms and related research areas. We tweaked our search strategy to address these challenges and avoid a result set of tens of thousands of potential studies.
A vital role was played by the IC & EC, which we applied twice. After the initial search, the first application was made to remove unrelated contributions that would introduce overheating into the snowballing. The second time was again after the snowballing to filter out wrongly selected contributions.